OpenStudy (anonymous):

how to solve lim (x-1/2x)/(2x+1/6x) x→∞

OpenStudy (anonymous):

\[ \large \displaystyle \lim_{x \to \infty} \frac{x-\frac{1}{2x}}{2x+\frac{1}{6x}} \] is that your question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well the leading terms in the nominator are x and 2x, does that already ring a bell? If not, have you tried factoring out x from both the nominator and the denominator?

OpenStudy (anonymous):

i got (1/2x)/{(12x^2+1)/6x}

OpenStudy (anonymous):

but i dont think i did it right....

OpenStudy (anonymous):

you can do it that way, find the common denominator and then turn out with your algebraic notation, what you think about this? \[ \large \displaystyle \lim_{x \to \infty} \frac{x-\frac{1}{2x}}{2x+\frac{1}{6x}}=\lim_{x \to \infty} \frac{x \left(1- \frac{2}{x^2} \right)}{x\left( 2+ \frac{1}{6x^2}\right)}=\lim_{x \to \infty} \frac{ \left(1- \frac{2}{x^2} \right)}{\left( 2+ \frac{1}{6x^2}\right)}\]

OpenStudy (anonymous):

how did you get \[\frac{ 2 }{ x^2 }\] when you factor out x?

OpenStudy (anonymous):

I factored out an \(x\) see the middle expression, you can see that they are equivalent if you distribute the \(x\) inside the parenthesis (...) again

OpenStudy (anonymous):

Oh okay! Thank you so much! :D

OpenStudy (anonymous):

you're very welcome, when solving such exercises it's also a good idea to get into the idea of 'what is the strongest term in the nominator/denominator', meaning which expressions grows fastest/largest or slowest

OpenStudy (anonymous):

I'll work with that idea. Thank you